# Breaking the Circle: An Autonomous Control-Switching Strategy for Stable Orographic Soaring in MAVs

### v1

@@ I Introduction
The flight endurance of micro air vehicles (MAVs) significantly constrains operational capabilities, limiting the scope of missions they can perform [1, 2]. This limitation is not solely due to inherently short flight durations, but also because take-off and landing procedures typically demand substantial time, energy, effort, and space. One potential solution to this problem lies in the advancement of battery technology, which could lead to improved efficiency. However, progress in this area has been relatively slow [3, 4]. Consequently, researchers have been exploring alternative solutions, such as using energy sources with higher energy densities or enabling mid-air refueling or recharging [5, 6]. Nevertheless, these approaches require considerable investment in hardware and system infrastructure, and often necessitate larger, heavier platforms—undermining the fundamental advantage of MAVs being small.
@@
An alternative approach is to exploit soaring, a flight technique widely employed by birds [7, 8, 9] and human-piloted glider aircraft [10, 11]. Soaring takes advantage of wind energy, specifically upward vertical winds, to gain altitude or remain airborne with minimal energy expenditure. A key strength of soaring is its compatibility with existing systems: it can be integrated into any fixed-wing aircraft without requiring hardware modifications, making it a valuable complement to other endurance-enhancing strategies.
@@
Various types of soaring techniques exist [12]. Dynamic soaring is a technique that capitalizes on the wind’s velocity gradient to gain energy, allowing migratory seabirds like albatrosses to traverse vast distances with minimal energy use [13, 14]. This technique has also inspired research efforts to implement dynamic soaring in unmanned aerial vehicles (UAVs), aiming to extend their endurance by autonomously harvesting energy from wind gradients [15, 16, 17]. Thermal soaring, exploiting updrafts generated by unevenly heated ground surfaces, is utilized by migratory birds and cross-country sailplanes [18, 19, 20]. Thermal soaring has been extensively studied in the context of control and guidance systems [21, 22, 23], as well as machine learning methods for autonomous MAVs [24, 25].
@@
Figure 1: Wind tunnel test trajectories of the Seal plane over 10 seconds using SAOS (proposed control switching method) and BASE (previous INDI-based controller from AOSoar [26] without any modifications) control methods. (a) SAOS: minimal circling and stable soaring flight. (b) BASE: pronounced circling behavior and trajectory divergence.
@@
Orographic soaring, another form of static soaring, utilizes updrafts created by natural or man-made obstacles such as dunes, mountains, or buildings [27, 28]. The Kestrel is a well-known example of a bird that employs this technique for hunting. Kestrels are often observed hovering in place over hills without flapping their wings as they search for prey—a behavior also referred to as wind-hovering [29, 30].
@@
Orographic soaring has not been widely studied. Early research focused on analyzing wind fields, flight conditions, and the feasibility of soaring, including in urban environments using simulations and wind measurements [31, 32, 33, 34]. Real-world demonstrations with MAVs were performed, although they were limited by reliance on prior wind field knowledge and predefined trajectories [35], or manual positioning before activating the soaring controller [36]. In wind tunnel environments, Suys et al. demonstrated autonomous orographic soaring with a glider MAV using a pitch controller [37]. While it required minimal wind field knowledge, some manual setup—such as setting a target gradient line—was still needed. Following that, AOSOAR achieved the first fully autonomous wind-hovering flight demonstration, requiring no human intervention, prior wind field knowledge, or precise initial positioning, using a powered fixed-wing MAV [26]. Utilizing an Incremental Nonlinear Dynamic Inversion (INDI) controller [38, 39] and an autonomous soaring position search method, it successfully performed wind-hovering flight with 0% throttle for over 30 minutes continuously. The method demonstrated the ability to adapt to changing wind conditions without any prior wind field data in the wind tunnel environment.
Although these studies have demonstrated autonomous soaring and wind-hovering, a specific flight pattern—circling—has been consistently observed. When circling remains within a small range, the MAV can still conserve more energy than during a typical flight, though it incurs additional energy costs due to actuator activity and throttle adjustments. In more severe cases, the circling radius expands, causing the MAV to drift out of the updraft or into no-wind zones, which presents a significant challenge. In wind tunnel environments, circling behavior could be reduced by controlling specific variables. Previous studies [37, 26] showed that circling was reduced under certain combinations of wind speed and slope angle. However, despite these improvements, circling behavior often still persisted. Moreover, in outdoor flights, manipulating the wind field is generally not feasible. Orographic soaring typically requires high wind speeds, which are often accompanied by strong gusts. These unpredictable conditions tend to worsen circling behavior and make stable soaring more difficult.
@@
3. Control mode switching: A control strategy is developed to transition the MAV from an underactuated to a fully actuated control mode during soaring, particularly when throttle input is near zero.
@@ II Circling in Orographic Soaring
Before delving into the proposed solutions to the circling behavior in orographic soaring, we first explain this problem in more detail. Circling is a flight behavior observed in orographic soaring, where the MAV fails to converge to the intended soaring position and instead overshoots it, resulting in continuous circular or elliptical flight patterns around the target location. This behavior typically arises when the MAV is near a feasible soaring region and consequently reduces its throttle to nearly zero. The issue primarily arises from the inherent characteristics of the orographic wind field, where both vertical and horizontal wind speeds vary along the slope, making soaring positions naturally unstable. Thus, it is not specific to any particular controller; however, the controller can influence the behavior, either mitigating or exacerbating it. To better understand this phenomenon, we present a comprehensive example that illustrates circling behavior in the context of orographic soaring.
@@
This behavior is more common in indoor environments than outdoors, such as the TU Delft open-jet facility (OJF) wind tunnel, where the limited height of the wind section tightly constrains vertical motion. To ensure precise altitude control under these constraints and reduce the risk of overshooting, relatively low altitude gains were used. In contrast, the longitudinal space is relatively unrestricted and poses less risk when overshooting, allowing for greater exploratory motion along the longitudinal axis. Consequently, longitudinal control gains were set relatively higher.
@@
The circling behavior occurs when the MAV attempts to approach a feasible soaring position, where the wind speed counteracts its airspeed and sink rate. However, due to its momentum, the MAV may slightly overshoot the reference position. As the MAV moves forward, both horizontal and vertical wind speeds decrease, resulting in additional forward and downward acceleration. To counteract this, the controller commands a pitch-up maneuver, inducing both backward and upward acceleration. This response leads to a diagonal upward trajectory rather than a direct return to the reference point. It is worth noting that the longitudinal and vertical axes respond at different rates to pitch adjustments, making error correction challenging—especially in non-uniform wind fields such as updrafts. When the MAV moves backward, the horizontal and vertical wind speeds increase, further accelerating it in the backward and upward directions. Upon reaching the reference position again, the MAV must cancel out the accumulated acceleration. However, with only the elevator available for control because the throttle is already zero while flying backward, achieving this correction instantaneously is difficult, leading to another overshoot. As the MAV moves toward the rear side of the slope, where the updraft is stronger, it gains altitude. In this case, horizontal acceleration can be corrected using throttle input. As a result, horizontal position errors are quickly mitigated, but altitude errors persist. To reduce altitude error, the MAV pitches down; however, this induces forward acceleration, leading to yet another overshoot. Unlike in backward flight, where the throttle can be used to counteract motion, forward acceleration is more difficult to control. Consequently, the MAV overshoots the reference position again and the cycle repeats, resulting in a persistent circling pattern.
@@
The other circling case, shown in figure 3, occurs more frequently in outdoor environments. In such settings, controller configurations typically differ from those used indoors due to the availability of larger airspace and differing flight scenarios. As a result, control gains are generally set higher for outdoor flights. For practical reasons, in outdoor operations, autonomous flights often begin from a higher altitude, and the vertical airspace tends to be less constrained than the horizontal space, which may be limited by obstacles or terrain features that generate updrafts. Accordingly, altitude gains are usually set relatively higher than longitudinal gains, which are set more conservatively. In this scenario, the controller responds more aggressively to altitude deviations, resulting in larger variations in pitch angle as it attempts to minimize vertical error. When the MAV overshoots the reference altitude and climbs too high, the controller commands a pitch-down maneuver to descend, causing the MAV to descend and move forward. As it descends, it overshoots the target altitude again, this time going too low. The controller then commands a pitch-up response to regain altitude, which drives the MAV upward and backward. This pattern repeats, leading to a persistent oscillatory behavior. Although the dynamics may look different from the indoor circling case, the underlying pattern remains the same: the MAV continuously overshoots the reference position due to aggressive gain settings and the coupling between pitch and longitudinal motion.
@@ III METHODOLOGY
We introduce improvements to the soaring controller to prevent or reduce circling behavior, using the controller from AOSOAR[26] as a baseline. AOSOAR has two main parts: the autonomous soaring search and the orographic soaring controller. We focused on improving the control part in this paper. For the control part of AOSOAR, an INDI-based controller with a weighted mean square (WMS) control allocation was used. It enables the MAV to move forward or backward in the updraft, and also cut off the throttle whenever required without changing any settings in-flight.
@@ III-A Angle of Attack (AoA) sensor
We implemented an AoA sensor on our MAVs using a 3D-printed wind vane with a servo motor to measure the angle. The servo motor was repurposed as a pure encoder by removing the motor part, with the angle directly measured and transmitted to the autopilot system. Despite minimal friction, the weight of the wind vane and remaining friction imposed a minimum operational wind speed for the sensor, approximately 6.0m/sm/s. Below this speed, measurements were inaccurate. However, since the minimum wind speed for soaring in both MAVs exceeds this threshold, the limitation had no significant impact.
@@ III-B AoA regulation through pitch control
With higher altitude or pitch gains, the MAV can stall due to a high angle of attack in the updraft. It can worsen circling behavior, lead to divergence, or potentially cause a crash. Since the wind field is non-uniform and wind components change with position, preventing stall solely through pitch measurements is not feasible. To address this, we used the AoA measurement to regulate it by limiting the pitch angle. The maximum AoA was determined based on the stall angle estimated by experiments in the wind tunnel for each MAV platform.
@@ III-C Outer loop control effectiveness with additional terms
In AOSOAR[26], the drag term was omitted, and lift estimation was simplified, assuming zero flight path angle. However, updraft soaring introduces unique dynamics, such as backward flight and high angles of attack, requiring adjustments. To better model these conditions, we incorporated the drag term and utilized AoA for force calculations.
@@
Then, the control equation can be derived by taking partial derivatives by applying a Taylor expansion to the Equation 2.
@@
Here, we assumed $\partialderivative{\theta}{\alpha}=1$. The wind flow is considered constant at a given position, and the wind direction is assumed to remain unchanged over time. Lift and drag forces were estimated and modeled using flight log data. However, accurately capturing the stall region was not feasible using flight data alone. Although a stall is ideally avoided during flight, incorporating this region into the model allows the controller to autonomously manage stall recovery. In practice, the aircraft may exceed the critical angle of attack due to turbulence, gusts, or variations in the wind field. Therefore, a realistic representation of stall behavior is essential, as opposed to abruptly dropping the lift force to zero. To address this limitation, we estimated the stall region using a NACA 1410 airfoil model. The lift coefficient versus angle of attack is shown in figure 7. A base model was selected based on wing measurements and then adjusted through offsets and scaling to better fit the observed data. This approach enables the use of a standardized airfoil model with minimal parameter tuning, facilitating adaptation across different MAV platforms.
@@ III-D Control switching method
Another challenge arises when the MAV must control its position along three axes (longitudinal, lateral, and vertical) but has only two available actuators, an elevator for pitch and ailerons for roll, particularly in the absence of throttle input. This situation typically occurs when the MAV is flying backward or near the feasible soaring position. The feasible soaring position refers to a location where the MAV can, in theory, perform wind-hovering, where the wind is sufficient to keep the MAV aloft without requiring thrust.
@@
To address this issue, we implemented a control-switching strategy. Instead of attempting to control all three axes simultaneously, the controller selectively manages either longitudinal or vertical position in conjunction with lateral position, ignoring one axis based on predefined settings. This approach prevents conflicts between the longitudinal and vertical axes when the throttle is already at 0% and pitch is the only available input for controlling both directions.
@@
When only altitude is controlled and the longitudinal position is left uncontrolled, the MAV responds to vertical disturbances autonomously. If the updraft is too strong, the MAV is naturally pushed upward. To correct this and maintain its target altitude, the MAV pitches down and begins to descend. If it then descends too much, it pitches up again to climb, which also causes it to move backward. Over time, this cycle continues until the MAV stabilizes in a region where the upward wind matches its natural sink rate—achieving an equilibrium that allows it to maintain position without active control in the longitudinal axis. A similar principle applies when the longitudinal position is controlled while altitude is left uncontrolled. If the updraft is too strong, the MAV is naturally pushed upward; if it is too weak, the MAV sinks. By actively maintaining its longitudinal position through control inputs, the MAV can passively adjust its altitude according to the available updraft. Over time, this allows the vehicle to settle at an equilibrium altitude—where the upward wind force balances its sink rate—without direct altitude control.
On the other hand, when the MAV is outside the soaring region or navigating toward a feasible soaring position, full three-axis control is required. Therefore, we applied a cascaded control allocation method that dynamically adjusts the control model in response to throttle saturation. The controller initially computes a standard three-axis control allocation. If the throttle becomes zero or is negatively saturated, a secondary control allocation is triggered using a reduced model that limits control to two degrees of freedom. Crucially, this approach does not permanently disable the throttle term in the controller; instead, it temporarily switches the outer-loop control logic as needed. The MAV retains the ability to re-engage throttle input when it is necessary. This strategy helps prevent the MAV from drifting too far backward beyond the reference position—a potentially unsafe condition—while preserving adaptability to wind gusts and environmental disturbances.

### v2

@@ I Introduction
The flight endurance of micro air vehicles (MAVs) significantly constrains operational capabilities, limiting the scope of missions they can perform [karydis2017energetics, bronz2009towards]. This limitation is not solely due to inherently short flight durations, but also because take-off and landing procedures typically demand substantial time, energy, effort, and space. One potential solution to this problem lies in the advancement of battery technology, which could lead to improved efficiency. However, progress in this area has been relatively slow [li2019practical, xu2023high]. Consequently, researchers have been exploring alternative solutions, such as using energy sources with higher energy densities or enabling mid-air refueling or recharging [de2021nederdrone, mohsan2022comprehensive]. Nevertheless, these approaches require considerable investment in hardware and system infrastructure, and often necessitate larger, heavier platforms—undermining the fundamental advantage of MAVs being small.
@@
An alternative approach is to exploit soaring, a flight technique widely employed by birds [pennycuick1972soaring, bildstein1987hunting, strandberg2006wind] and human-piloted glider aircraft [metzger1975optimal, akos2008comparing]. Soaring takes advantage of wind energy, specifically upward vertical winds, to gain altitude or remain airborne with minimal energy expenditure. A key strength of soaring is its compatibility with existing systems: it can be integrated into any fixed-wing aircraft without requiring hardware modifications, making it a valuable complement to other endurance-enhancing strategies.
@@
Various types of soaring techniques exist [mohamed2022opportunistic]. Dynamic soaring is a technique that capitalizes on the wind’s velocity gradient to gain energy, allowing migratory seabirds like albatrosses to traverse vast distances with minimal energy use [weimerskirch2000fast, sachs2012flying]. This technique has also inspired research efforts to implement dynamic soaring in unmanned aerial vehicles (UAVs), aiming to extend their endurance by autonomously harvesting energy from wind gradients [bonnin2015energy, mir2018review, park2025application]. Thermal soaring, exploiting updrafts generated by unevenly heated ground surfaces, is utilized by migratory birds and cross-country sailplanes [harel2016adult, akos2010thermal, keskin2025adaptive]. Thermal soaring has been extensively studied in the context of control and guidance systems [edwards2008implementation, cobano2013thermal, depenbusch2018autosoar], as well as machine learning methods for autonomous MAVs [reddy2018glider, flato2024revealing].
@@
Figure 1: Wind tunnel test trajectories of the Seal plane over 10 seconds using SAOS (proposed control switching method) and BASE (previous INDI-based controller from AOSoar [hwang2023aosoar] without any modifications) control methods. (a) SAOS: minimal circling and stable soaring flight. (b) BASE: pronounced circling behavior and trajectory divergence.
@@
Orographic soaring, another form of static soaring, utilizes updrafts created by natural or man-made obstacles such as dunes, mountains, or buildings [bohrer2012estimating, sage2019orographic]. The Kestrel is a well-known example of a bird that employs this technique for hunting. Kestrels are often observed hovering in place over hills without flapping their wings as they search for prey—a behavior also referred to as wind-hovering [videler1983intermittent, penn2022method].
@@
Orographic soaring has not been widely studied. Early research focused on analyzing wind fields, flight conditions, and the feasibility of soaring, including in urban environments using simulations and wind measurements [white2012soaring, white2012feasibility, guerra2020unmanned, lu2025energy]. Real-world demonstrations with MAVs were performed, although they were limited by reliance on prior wind field knowledge and predefined trajectories [fisher2015emulating], or manual positioning before activating the soaring controller [de2021never]. In wind tunnel environments, Suys et al. demonstrated autonomous orographic soaring with a glider MAV using a pitch controller [suys2023autonomous]. While it required minimal wind-field knowledge, some manual setup—such as setting a target gradient line—was still needed. Following that, AOSOAR achieved the first fully autonomous wind-hovering flight demonstration, requiring no human intervention, prior wind field knowledge, nor precise initial positioning, using a powered fixed-wing MAV [hwang2023aosoar]. Utilizing an Incremental Nonlinear Dynamic Inversion (INDI) controller [sieberling2010robust, smeur2020incremental] and an autonomous soaring position search method, it successfully performed wind-hovering flight with 0% throttle for over 30 minutes continuously. The method demonstrated the ability to adapt to changing wind conditions without any prior wind-field data in the wind tunnel environment.
Although these studies have demonstrated autonomous soaring and wind-hovering, a specific flight pattern—circling—has been consistently observed. When the amplitude of the circling remains small, the MAV can still harvest a lot of energy from the updraft, though it incurs additional energy losses due to actuator activity and throttle adjustments. In more severe cases, the circling radius expands, causing the MAV to drift out of the updraft or into no-wind zones, which presents a significant challenge. In wind tunnel environments, circling behavior could be reduced by controlling specific variables. Previous studies [suys2023autonomous, hwang2023aosoar] showed that circling was reduced under certain combinations of wind speed and slope angle. However, despite these improvements, circling behavior often persisted. Moreover, in outdoor flights, manipulating the wind field is generally not feasible. Orographic soaring typically requires high wind speeds, which are often accompanied by strong gusts. These unpredictable conditions tend to worsen circling behavior and make stable soaring more difficult.
@@
3. Control-mode switching: A control strategy is developed to transition the MAV from an underactuated to a fully actuated control mode during soaring, particularly when the throttle input is near zero.
@@ II Circling in Orographic Soaring
Before delving into the proposed solutions to the circling behavior in orographic soaring, we first explain this problem in more detail. Circling is a flight behavior observed in orographic soaring, where the MAV fails to converge to the intended soaring position but instead continuously overshoots it in the perpendicular axis, resulting in continuous circular or elliptical flight patterns around the target location. The issue primarily arises from the inherent characteristics of the orographic wind field, where both vertical and horizontal wind speeds vary along the slope. A perturbation in one axis, therefore, also creates a perturbation in the other axis, which reduces the stability of soaring at a fixed position. While it is not specific to any particular controller, the controller can influence the behavior, either mitigating or exacerbating it. To better understand this phenomenon, we present a comprehensive example that illustrates circling behavior in the context of orographic soaring.
@@
This behavior is more common in indoor environments than outdoors, such as the open-jet facility (OJF) wind tunnel, where the limited height of the wind section tightly constrains vertical motion. To avoid aggressive oscillations or overshooting under these tight constraints, relatively low altitude control gains were used, allowing smoother but slower vertical responses. In contrast, the longitudinal axis is relatively less constrained and more tolerant of overshooting, so higher longitudinal control gains were employed to better regulate forward and backward motion.
@@
The circling behavior occurs when the MAV attempts to approach a feasible soaring position, where the wind speed counteracts its airspeed and sink rate. However, due to its momentum, the MAV may slightly overshoot the reference position. When, for instance, the MAV overshoots in the forward direction, it arrives in a section of the updraft with lower horizontal and vertical wind speeds, resulting in additional forward and downward acceleration. The forward overshoot then induces a small error in altitude. To counteract this, the controller commands a pitch-up maneuver, inducing both backward and upward acceleration. This response leads to a diagonal upward trajectory rather than a direct return to the reference point. It is worth noting that the longitudinal and vertical axes respond at different rates to pitch adjustments, making error correction challenging—especially in non-uniform wind fields such as updrafts. When the MAV moves backward, it enters areas with stronger horizontal and vertical wind speeds, further accelerating it in the backward and upward directions. Upon reaching the reference position again, the MAV must cancel out the accumulated acceleration. However, with only the elevator available for control because the throttle is already zero while flying backward, achieving this correction instantaneously is difficult, possibly leading to another overshoot. As the MAV moves toward the rear side of the slope, where the updraft is stronger, it gains altitude. In this case, horizontal acceleration can also be corrected using throttle input. As a result, horizontal position errors are quickly mitigated, but altitude errors persist. To reduce altitude error, the MAV pitches down; however, this induces forward acceleration, leading to yet another overshoot. Unlike in backward flight, where the throttle can be used to counteract motion, forward acceleration is more difficult to control. Consequently, the MAV overshoots the reference position again and the cycle repeats, resulting in a persistent circling pattern.
The other circling case, shown in figure 3, occurs more frequently in outdoor environments. In such settings, controller configurations typically differ from those used indoors due to the availability of larger airspace and differing flight scenarios.
@@
For practical reasons, autonomous flights in outdoor conditions often begin from higher altitudes, and the vertical airspace is typically less constrained than the horizontal space, which may be limited by terrain or obstacles generating updrafts. Accordingly, altitude control gains are often tuned more aggressively than longitudinal ones to maintain altitude in a wide vertical range. However, this increased sensitivity can result in more reactive pitch behavior. When the MAV overshoots the reference altitude and climbs too high, the controller pitches down to descend, causing forward motion. After descending, it may overshoot again and go too low, prompting a pitch-up response that sends it backward and upward. This coupling between altitude and longitudinal motion leads to a repeating oscillatory pattern. Although this behavior differs visually from the circling observed indoors, it stems from the same underlying issue: coupled dynamics between pitch and acceleration in a non-uniform wind field.
@@ III METHODOLOGY
We introduce improvements to the soaring controller to prevent or reduce circling behavior, using the controller from AOSOAR[hwang2023aosoar] as a baseline. AOSOAR has two main parts: the autonomous soaring search and the orographic soaring controller. We focused on improving the control part in this paper. For the control part of AOSOAR, an INDI-based controller with a weighted mean square (WMS) control allocation was used. It enables the MAV to move forward or backward in the updraft, and also cut off the throttle whenever required without changing any settings in-flight.
@@ III-A Angle of Attack (AoA) sensor
We implemented an AoA sensor on our MAVs using a 3D-printed wind vane with a potentiometer to measure the angle. This was done by repurposing a servo actuator as a pure encoder by removing the electric motor, while the angle is still directly measured and transmitted to the autopilot system. Despite minimal friction, the weight of the wind vane and remaining friction imposed a minimum operational wind speed for the sensor, approximately 6.0m/sm/s. Below this speed, measurements were inaccurate. However, since the minimum wind speed for soaring in both MAVs exceeds this threshold, the limitation had no significant impact.
@@ III-B AoA regulation through pitch control
With higher altitude or pitch gains, the MAV can stall due to a high angle of attack in the updraft. It can worsen the circling behavior, lead to divergence, or potentially cause a crash. Since the wind field is non-uniform and wind components change with position, preventing stall solely through pitch measurements is not feasible. To address this, we used the AoA measurement to regulate it by limiting the pitch angle. The maximum AoA was determined based on the stall angle estimated by experiments in the wind tunnel for each MAV platform, using:
@@ III-C Outer loop control effectiveness with additional terms
In AOSOAR[hwang2023aosoar], the drag term was omitted, and lift estimation was simplified, assuming zero flight path angle. However, updraft soaring introduces unique dynamics, such as backward flight and high angles of attack, requiring adjustments. To better model these conditions, we incorporated the drag term and utilized AoA for force calculations.
@@
Then, the control equation can be derived by taking partial derivatives by applying a Taylor expansion to the Equation 2:
@@
Here, we assumed $\partialderivative{\theta}{\alpha}=1$. The wind flow is considered constant at a given position, and the wind direction is assumed to remain unchanged over time. Lift and drag forces were estimated and modeled using flight-log data. However, accurately capturing the stall region was not feasible using flight data alone. Although a stall is ideally avoided during flight, incorporating this region into the model allows the controller to autonomously manage stall recovery. In practice, the aircraft may exceed the critical angle of attack due to turbulence, gusts, or variations in the wind field. Therefore, a realistic representation of stall behavior is essential, as opposed to abruptly dropping the lift force to zero. To address this limitation, we estimated the stall region using a NACA 1410 airfoil model. The lift coefficient versus angle of attack is shown in figure 7. A base model was selected based on wing measurements and then adjusted through offsets and scaling to better fit the observed data. This approach enables the use of a standardized airfoil model with minimal parameter tuning, facilitating adaptation across different MAV platforms.
@@ III-D Control switching method
Another challenge arises when the MAV must control its position along three axes (longitudinal, lateral, and vertical) but has only two available actuators, an elevator for pitch and ailerons for roll, particularly in the absence of throttle input. This situation typically occurs when the MAV has too much total energy near the feasible soaring position. The feasible soaring position refers to a location where the MAV can, in theory, perform wind-hovering, where the wind is sufficient to keep the MAV aloft without requiring thrust.
@@
To address this issue, we implemented a control-switching strategy. Instead of attempting to control all three axes simultaneously, the controller selectively manages either longitudinal or vertical position in conjunction with lateral position, ignoring one axis based on predefined settings. This approach solves the underactuation problem between the longitudinal and vertical axes when the throttle is already at 0% and pitch is the only available input for controlling both directions.
@@
When only altitude is controlled and the longitudinal position is left uncontrolled, the MAV responds to vertical disturbances autonomously. If the updraft is too strong, the MAV is naturally pushed upward. To correct this and maintain its target altitude, the MAV pitches down and begins to move forward to a region with weaker updrafts. If too low, the resulting pitch correction will make the MAV fly slower, moving backwards towards regions with stronger updrafts. Over time, this cycle continues until the MAV stabilizes in a region where the upward wind matches its natural sink rate—achieving an equilibrium that allows it to maintain position without active control in the longitudinal axis. A similar principle applies when the longitudinal position is controlled while the altitude is left uncontrolled. If the updraft is too strong, the MAV is naturally pushed upward; if it is too weak, the MAV sinks. By actively maintaining its longitudinal position through control inputs, the MAV can passively adjust its altitude according to the available updraft. This is made possible by the natural tendency of updrafts to decrease with altitude. Over time, this allows the vehicle to settle at an equilibrium altitude—where the upward wind force balances its sink rate—without direct altitude control.
On the other hand, when the MAV is outside the soaring region or navigating toward a feasible soaring position, full three-axis control is required. Therefore, we applied a cascaded control allocation method that dynamically adjusts the control model in response to throttle saturation. The controller initially computes a standard three-axis control allocation. If the throttle becomes zero or is negatively saturated, a secondary control allocation is triggered using a reduced model that limits control to two degrees of freedom. Crucially, this approach does not permanently disable the throttle term in the controller; instead, it temporarily switches the outer-loop control logic as needed. The MAV retains the ability to re-engage throttle input whenever it is necessary. This strategy helps prevent the MAV from drifting too far backward beyond the reference position—a potentially unsafe condition—while preserving adaptability to wind gusts and environmental disturbances.